OpenStudy (anonymous):

For the vectors v1=(1,0,2,2) v2=(1,1,0,3), v3=(1,4,0,0) find a linear homogeneous system which has the span of these vectors as solution space.

OpenStudy (anonymous):

@ivanmlerner do you know how to do this one?

OpenStudy (anonymous):

Just a second, I need to look some words in the dictionary, I'm not familiar with this subject in english.

OpenStudy (anonymous):

Look, I have a way, but since I haven't had linear algebra beyond the most basics this is probably not the best approach. We know that the solutions of a homogeneous linear system are given by av1+bv2+cv3 And that this linear system must have the following form: a1x+b1y+c1z+d1h=0 a2x+b2y+c2z+d2h=0 a3x+b3y+c3z+d3h=0 a4x+b4y+c4z+d4h=0 Lets take the first of these equations, now we have 3 points that must satisfy this equation and that are linearly independent(not sure if that is the term you use), and we can get a fourth just by making a linear combination of these 3. Doing this, we get an equation with 4 unknown constants and four points that mus satisfy this equation, therefore we have a 4x4 system with a1, b1, c1, d1 as the solution. Doing that for all of the four equations, we find all the constants needed.

OpenStudy (anonymous):

so are you saying that the system we are finding is one equation, or 4 equations?

OpenStudy (anonymous):

4 equations to find the 4 constants of the first equation of the original system, and the same thing for the other 3

OpenStudy (anonymous):

i was thinking that you set those vectors in a matrix and you do reduced echelon form, but didnt know what to do after that... i am not sure if i quite understand how to make the four equations .... what are the four equations i would use for (1,0,2,2)

OpenStudy (anonymous):

I don't know what this method that you said is, because as I said I haven

OpenStudy (anonymous):

what course did you take that you learned some of this stuff from?

OpenStudy (anonymous):

t studied that so far, but what I was thinking was not make 4 equations for each poin, but make for equations using all points for each equation of the system

OpenStudy (anonymous):

Highschool, just some simple methods for solving linear systems and nothing greater than 3 dimensions, but even though I don't have any of the other methods, mine usually work, even though sometimes they are not very eficient

OpenStudy (anonymous):

well if they work, then id like to learn how to use them, but i don't think I quite understand how to do your method.

OpenStudy (anonymous):

the system you want is this: a1x+b1y+c1z+d1h=0 a2x+b2y+c2z+d2h=0 a3x+b3y+c3z+d3h=0 a4x+b4y+c4z+d4h=0 Looking just at the first equation and suposing all points and its linear combinations are solutions of it: a1+2c1+2d1=0 (point (1, 0, 2, 2)) a1+b1+3d1=0 (point (1, 1, 0, 3)) a1+4b1=0 (point (1, 4, 0, 0)) 2a1+b1+2c1+5d1=0 (Point v1+v2=(2, 1, 2, 5)) Since the span of these vectors must be a solution v1+v2=(2, 1, 2, 5) must also be a solution. Now you have a linear system wich solution gives you a1, b1, c1, d1. This system so not have only one solution, but thats ok, since you want a system that gives you some restrictions so that you can create a system that has the span os these points as a solution. To sumarize, your problem has many valid solutions.

OpenStudy (anonymous):

is there a reason we have to have 4 equations for our system? why cant we have just one?

OpenStudy (anonymous):

Well, we have 3 vectors that are independent, and all of them must satisfy the equation so I was wrong to "create" a point to make a fouth equation, but you need to have at least 3, because otherwise you are not aplying all the conditions and the equation might not satisfy all the points, giving a wrong answer.

OpenStudy (anonymous):

If I gave you a vector, lets call it "u"... and if i was to multiply this vector by v1, v2, and v3 such that v1u=v2u=v3u=0..... would that satisfy what I am looking for or no??

OpenStudy (anonymous):

You mean a vector that is perpendicular to all the three? I thought about it too because the equation is similar to the equation of a plane, but I got consused because of the 4 dimensions, so I just followed another way, but maybe that is true.

OpenStudy (anonymous):

yeah im not really sure either... i do thank you for your help though. it sounds like you understand this stuff better than i do, and im in the class.

OpenStudy (anonymous):

you are welcome, the methods you are learning are much more powerfull though, let me know if you get the result with this method

OpenStudy (anonymous):